# Beyond Characteristic Equations: A Unified Non-Bloch Band Theory via Wavefunction Data

### v1

@@ Abstract
We test this conclusion using a non-Hermitian Hamiltonian belonging to the symplectic class. The Bloch Hamiltonian is given by $H(k)=\begin{pmatrix}0&D_{1}(k)\\ D_{2}(k)&0\end{pmatrix}$ where $D_{1}(k)=t\sin k\,\sigma_{x}+\left(\Delta+u+u\cos k+i\frac{\gamma}{2}\right)\sigma_{y}$, $D_{2}(k)=t\sin k\,\sigma_{x}+\left(\Delta+u+u\cos k\right)\sigma_{y}$, and $\sigma_{i}$ are Pauli matrices. This system has a symplectic TRS operator $\mathcal{U}_{T}=-i\sigma_{y}\otimes\mathbb{I}_{2\times 2}$, under which the Hamiltonian satisfies $\mathcal{U}_{T}H(k)^{T}\mathcal{U}_{T}^{-1}=H(-k)$. In Fig. 3, we plot the end-to-end exponential factor $\alpha_{\leftarrow}$ and compare it with the solutions of the characteristic equation $\text{det}[\omega-H(\beta)]=0$. This equation yields eight analytical solutions, which we label using Roman numerals for clarity. Among them, four are shown in Fig. 3(a); the other four consist of two solutions ($\beta_{\text{IV}},\beta_{\text{VII}}$) that lie above and two ($\beta_{\text{III}},\beta_{\text{VIII}}$) that lie below those depicted. As shown in Fig. 3(a), there are 4 crossing points, which are potential transition points of the amplification factor $\alpha_{\leftarrow}$. Based on these crossings, the figure can be divided into four distinct regions, labeled A, B, C and D. As an illustrative example, we focus on region B. In this region, the completeness relation $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ vanishes for the index set $P=\{\text{III},\text{VIII},\text{I},\text{V}\}$, while it remains nonzero for $P=\{\text{III},\text{VIII},\text{I},\text{VI}\}$. These results are obtained from the structure of the wavefunctions (see SM [40]). Therefore, according to the criterion in Eq. (9), the minimal index set is $P_{\text{min}}=\{\text{III},\text{VIII},\text{I},\text{VI}\}$, and the corresponding spatial exponential factor $\alpha_{\leftarrow}=|\beta_{\text{VI}}|$.
@@
Figure 3: Comparison between the solutions of the characteristic equation $\text{det}[\omega-H(\beta)]=0$ and the end-to-end exponential factor $\alpha_{\leftarrow}$. Parameters are set as $t=1.01,u=1,\gamma=1.2$. (a) Magnitudes of analytical solutions $\beta_{i}$ of the characteristic equation plotted as a function of the parameter $\Delta$. (b) End-to-end exponential factor $\alpha_{\leftarrow}$ as a function of the parameter $\Delta$, obtained via a linear fit of the logarithm of the end-to-end Green’s function versus the system length $L$, with $L\in[150,170]$.
@@
A universal feature is also observed in Fig. 3(b): the spatial exponential factor $\alpha_{\leftarrow}$ exhibits a turning point only at the boundary between regions B and region C, while the crossing points at $|\beta|=1$ (or $\text{log}(|\beta|)=0$) are not associated with such transitions. This is because the crossing at $|\beta|=1$ violates the completeness condition due to an exchange of wavefunctions, a result that can be rigorously proven in SM [40]. In summary, according to Eq. (9) and Eq. (10), we find that $\alpha_{\leftarrow}=|\beta_{\text{VI}}|$ in regions A and B, and $\alpha_{\leftarrow}=|\beta_{\text{II}}|$ in regions C and D. This prediction is in excellent agreement with the numerical results shown in Fig. 3(b). However, the question remains: is $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ or $\alpha_{\leftarrow}=|\beta_{M+1}(\omega)|$? Since $|\beta_{M}(\omega)|$ and $|\beta_{M+1}(\omega)|$ are defined by their magnitudes in ascending order, we can immediately deduce from Fig. 3(a) that $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ in regions A and D and $\alpha_{\leftarrow}=|\beta_{M+1}(\omega)|$ in regions B and C. This stands in stark contrast to the single-band case, where $\alpha_{\leftarrow}$ always equals to $|\beta_{M}(\omega)|$. In symplectic class systems, $\alpha_{\leftarrow}$ can take either value, and its precise value is determined by Eq. (9) and Eq. (10). Importantly, distinguishing whether $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ or $|\beta_{M+1}(\omega)|$ is crucial for predicting signal amplification, because $|\beta_{M+1}(\omega)|$ which is larger than 1 stands for signal amplification while $|\beta_{M}(\omega)|$ which is smaller than 1 stands for signal decay.
@@
This behavior of the spatial exponential factor $\alpha_{\leftarrow}$ also reveals a deeper connection to the GBZ condition. Specifically, a turning point in $\alpha_{\leftarrow}$ indicates that the resonance frequency $\omega$ enters the spectrum. The universal feature we observed, namely that crossing points at $|\beta|=1$ do not correspond to turning points, implies that the GBZ condition in the symplectic class cannot take the conventional form $|\beta_{M}|=|\beta_{M+1}|$. Interestingly, some previous works have proposed an even stronger and more definitive statement: that the GBZ condition is given by $|\beta_{M+1}|=|\beta_{M+2}|$. However, we emphasize that the validity of this conclusion is contingent upon a prerequisite: there are no other symmetries further alter the completeness among the wavefunctions. For example, if the system possesses two distinct symplectic symmetries $\mathcal{U}_{T,1}$ and $\mathcal{U}_{T,2}$, then the spatial exponential factor $\alpha_{\leftarrow}$ may instead correspond to $|\beta_{M+2}(\omega)|$, and the GBZ condition may accordingly shift away from $|\beta_{M+1}|=|\beta_{M+2}|$. Nevertheless, regardless of how many symmetries the system possesses, our expressions for the spatial exponential factor $\alpha_{\leftarrow}$ in Eq. (9) and Eq. (10) remain universally valid. These formulas provide a symmetry-agnostic method for determining $\alpha_{\leftarrow}$, thereby offering a consistent framework even in the presence of complex symmetry structures.
@@
Discussion and conclusion.— The results presented in this work focus on one-dimensional non-Hermitian systems. For systems in higher dimensions, the Amoeba formulation has been developed as a nontrivial generalization of one-dimensional non-Bloch band theory [26]. It is worth noting, however, that the Amoeba formulation still relies solely on the characteristic equation as input, and the issue of GBZ condition shifting remains unresolved [36]. We believe that certain aspects of our results for one-dimensional systems may still be applicable to higher-dimensional cases. For instance, in a two-dimensional system with Hamiltonian $H(\beta_{x},\beta_{y})$, where $\beta_{x}$ and $\beta_{y}$ are the generalized complex momenta, one may fix $\beta_{y}$ at a value corresponding to a point on the two-dimensional GBZ (i.e., a minimal point in the Amoeba) and apply our theory to the resulting pseudo-one-dimensional system along the $x$-direction. In such cases, we expect that the biorthogonal completeness condition given in Eq. (7) remains a necessary condition. Whether Eq. (7) also serves as a sufficient condition in higher dimensions, and how to more effectively treat such systems, are important questions that we leave for future study.
@@
Acknowledgements.—This work is supported by the China Postdoctoral Science Foundation 2023M743398.

### v2

@@ Abstract
We test this conclusion using a non-Hermitian Hamiltonian belonging to the symplectic class. The Bloch Hamiltonian is given by $H(k)=\begin{pmatrix}0&D_{1}(k)\\ D_{2}(k)&0\end{pmatrix}$ where $D_{1}(k)=t\sin k\,\sigma_{x}+\left(\Delta+u+u\cos k+i\frac{\gamma}{2}\right)\sigma_{y}$, $D_{2}(k)=t\sin k\,\sigma_{x}+\left(\Delta+u+u\cos k\right)\sigma_{y}$, and $\sigma_{i}$ are Pauli matrices. This system has a symplectic TRS operator $\mathcal{U}_{T}=-i\sigma_{y}\otimes\mathbb{I}_{2\times 2}$, under which the Hamiltonian satisfies $\mathcal{U}_{T}H(k)^{T}\mathcal{U}_{T}^{-1}=H(-k)$. In Fig. 3, we plot the end-to-end exponential factor $\alpha_{\leftarrow}$ and compare it with the solutions of the characteristic equation $\text{det}[\omega-H(\beta)]=0$. This equation yields eight analytical solutions, which we label using Roman numerals for clarity. Among them, four are shown in Fig. 3(a); the other four consist of two solutions ($\beta_{\text{IV}},\beta_{\text{VII}}$) that lie above and two ($\beta_{\text{III}},\beta_{\text{VIII}}$) that lie below those depicted. As shown in Fig. 3(a), there are four crossing points, which are potential transition points of the amplification factor $\alpha_{\leftarrow}$. Based on these crossings, the figure can be divided into four distinct regions, labeled A, B, C, and D. As an illustrative example, we focus on region B. In this region, the completeness relation $\text{det}\left(\sum_{i\in P}x_{i}y_{i}^{T}\right)$ vanishes for the index set $P=\{\text{III},\text{VIII},\text{I},\text{V}\}$, while it remains nonzero for $P=\{\text{III},\text{VIII},\text{I},\text{VI}\}$. These results are obtained from the structure of the wavefunctions (see the Supplemental Material [50]). Therefore, according to the criterion in Eq. (9), the minimal index set is $P_{\text{min}}=\{\text{III},\text{VIII},\text{I},\text{VI}\}$, and the corresponding spatial exponential factor $\alpha_{\leftarrow}=|\beta_{\text{VI}}|$.
@@
Figure 3: Comparison between the solutions of the characteristic equation $\text{det}[\omega-H(\beta)]=0$ and the end-to-end exponential factor $\alpha_{\leftarrow}$. Parameters are set as $t=1.01,u=1,\gamma=1.2$, and $\omega=0.01i$. (a) Magnitudes of analytical solutions $\beta_{i}$ of the characteristic equation plotted as a function of the parameter $\Delta$. (b) End-to-end exponential factor $\alpha_{\leftarrow}$ as a function of the parameter $\Delta$, obtained via a linear fit of the logarithm of the end-to-end Green’s function vs the system length $L$, with $L\in[150,170]$.
@@
A universal feature is also observed in Fig. 3(b): The spatial exponential factor $\alpha_{\leftarrow}$ exhibits a turning point only at the boundary between regions B and C, while the crossing points at $|\beta|=1$ (or $\text{log}(|\beta|)=0$) are not associated with such transitions. This is because the crossing at $|\beta|=1$ violates the completeness condition due to an exchange of wavefunctions, a result that can be rigorously proven in the Supplemental Material [50]. In summary, according to Eqs. (9) and (10), we find that $\alpha_{\leftarrow}=|\beta_{\text{VI}}|$ in regions A and B, and $\alpha_{\leftarrow}=|\beta_{\text{II}}|$ in regions C and D. This prediction is in excellent agreement with the numerical results shown in Fig. 3(b). However, the question remains: Is $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ or $\alpha_{\leftarrow}=|\beta_{M+1}(\omega)|$? Since $|\beta_{M}(\omega)|$ and $|\beta_{M+1}(\omega)|$ are defined by their magnitudes in ascending order, we can immediately deduce from Fig. 3(a) that $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ in regions A and D and $\alpha_{\leftarrow}=|\beta_{M+1}(\omega)|$ in regions B and C. This stands in stark contrast to the single-band case, where $\alpha_{\leftarrow}$ is always equal to $|\beta_{M}(\omega)|$. In symplectic class systems, $\alpha_{\leftarrow}$ can take either value, and its precise value is determined by Eqs. (9) and (10). Importantly, distinguishing whether $\alpha_{\leftarrow}=|\beta_{M}(\omega)|$ or $|\beta_{M+1}(\omega)|$ is crucial for predicting signal amplification, because $|\beta_{M+1}(\omega)|$ that is larger than 1 stands for signal amplification while $|\beta_{M}(\omega)|$ that is smaller than 1 stands for signal decay.
@@
This behavior of the spatial exponential factor $\alpha_{\leftarrow}$ also reveals a deeper connection to the GBZ condition. Specifically, a turning point in $\alpha_{\leftarrow}$ indicates that the resonance frequency $\omega$ enters the spectrum. The universal feature we observed, namely, that crossing points at $|\beta|=1$ do not correspond to turning points, implies that the GBZ condition in the symplectic class cannot take the conventional form $|\beta_{M}|=|\beta_{M+1}|$. Interestingly, some previous works have proposed an even stronger and more definitive statement that the GBZ condition is given by $|\beta_{M+1}|=|\beta_{M+2}|$. However, we emphasize that the validity of this conclusion is contingent upon a prerequisite: There are no other symmetries that further alter the completeness among the wavefunctions. For example, if the system possesses two distinct symplectic symmetries $\mathcal{U}_{T,1}$ and $\mathcal{U}_{T,2}$, then the spatial exponential factor $\alpha_{\leftarrow}$ may instead correspond to $|\beta_{M+2}(\omega)|$, and the GBZ condition may accordingly shift away from $|\beta_{M+1}|=|\beta_{M+2}|$. Nevertheless, regardless of how many symmetries the system possesses, our expressions for the spatial exponential factor $\alpha_{\leftarrow}$ in Eqs. (9) and (10) remain universally valid. These formulas provide a symmetry-agnostic method for determining $\alpha_{\leftarrow}$, thereby offering a consistent framework even in the presence of complex symmetry structures.
@@
Discussion and conclusion.— The results presented in this work focus on one-dimensional non-Hermitian systems. For systems in higher dimensions, the amoeba formulation has been developed as a nontrivial generalization of one-dimensional non-Bloch band theory [31]. It is worth noting, however, that the amoeba formulation still relies solely on the characteristic equation as input, and the issue of GBZ condition shifting remains unresolved [41]. We believe that certain aspects of our results for one-dimensional systems may still be applicable to higher-dimensional cases. For instance, in a two-dimensional system with Hamiltonian $H(\beta_{x},\beta_{y})$, where $\beta_{x}$ and $\beta_{y}$ are the generalized complex momenta, one may fix $\beta_{y}$ at a value corresponding to a point on the two-dimensional GBZ (i.e., a minimal point in the amoeba) and apply our theory to the resulting pseudo-one-dimensional system along the $x$-direction. In such cases, we expect that the biorthogonal completeness condition given in Eq. (7) remains a necessary condition. Whether Eq. (7) also serves as a sufficient condition in higher dimensions and how to more effectively treat such systems are important questions that we leave for future study.
The one-dimensional non-Bloch band theory has already been verified in experimental settings [57]. However, the nonstandard GBZ conditions in highly symmetric non-Hermitian systems have not yet been experimentally demonstrated. Our modified theory not only provides a non-ad hoc framework at the theoretical level, but its generality also makes it well suited for systematic experimental validation of GBZ theory in such symmetric non-Hermitian systems. Experimentally, mechanical oscillators [58, 59], photonic crystals [6, 7, 8], and electrical circuit [60, 61] represent promising platforms for such investigations.
@@
Acknowledgement.—This work was supported by the China Postdoctoral Science Foundation under Grant No. 2023M743398.